Determine the force between a point charge of -100 μc and +50 μc when they are 200 cm apart in a vacuum.


1
Expert's answer
2021-05-11T07:24:21-0400

According to coulomb's law ,


F= "k\times \frac{q_1q_2}{d^2}"


F= "9\times 10^9 \times \frac{100\times 10^{-6} \times 50 \times 10^{-6}}{2}" = 22.5 N

F= 22.5 N